# Towards More Realistic Extraction Attacks: An Adversarial Perspective

### v2

@@ Abstract
Language models are prone to memorizing parts of their training data which makes them vulnerable to extraction attacks. Existing research often examines isolated setups–such as evaluating extraction risks from a single model or with a fixed prompt design. However, a real-world adversary could access models across various sizes and checkpoints, as well as exploit prompt sensitivity, resulting in a considerably larger attack surface than previously studied. In this paper, we revisit extraction attacks from an adversarial perspective, focusing on how to leverage the brittleness of language models and the multi-faceted access to the underlying data. We find significant churn in extraction trends, i.e., even unintuitive changes to the prompt, or targeting smaller models and earlier checkpoints, can extract distinct information. By combining information from multiple attacks, our adversary is able to increase the extraction risks by up to 2×2$\times$2 ×. Furthermore, even with mitigation strategies like data deduplication, we find the same escalation of extraction risks against a real-world adversary. We conclude with a set of case studies, including detecting pre-training data, copyright violations, and extracting personally identifiable information, showing how our more realistic adversary can outperform existing adversaries in the literature.Equal contribution [^2] Code released at https://github.com/EQUAL-Mila/llm_extraction_eval
@@ 1 Introduction
Large language models (LLMs) have grown considerably in size (Meta AI, 2024; Zhao et al., 2023), and have become integral to a wide range of tasks such as knowledge retrieval, question answering, code generation, machine translation, etc.
To complement this growing scale, LLMs are often trained on large amounts of data (Penedo et al., 2024; Soboleva et al., 2023; Gao et al., 2020; Raffel et al., 2020) that may include private, unlicensed or copyrighted information, especially if directly scraped from the web. As LLMs are prone to memorizing the data they’ve been trained on, they can be prompted to expose sensitive contexts - making it easier for an adversary to extract information in a black-box setting. Naturally, a question arises, how big is the risk imposed due to memorization?
Extraction attacks offer an empirical framework to quantify the information leakage in the presence of an adversary. The most commonly studied extraction attack is discoverable memorization (Carlini et al., 2023; Kassem et al., 2024), where the model is prompted with a portion of a sentence from the training data to extract the rest, thus enabling the adversary to perform targeted attacks.
Current extraction attacks study memorization trends in LLMs across isolated settings like model sizes, generation hyperparameters and learning dynamics (Carlini et al., 2021). While effective, they underestimate the risk posed due to a multi-faceted access to the underlying data in the current LLM ecosystem. For instance, we show that an adversary can exploit the sensitivity of LLMs to prompt structure, length and content, to amplify the information gained. The current accessibility to frequently updated model sizes (Meta AI, 2024); checkpoints (Biderman et al., 2023b; Groeneveld et al., 2024) and a large array of model families such as Llama (Meta AI, 2024), Gemini  (Team et al., 2023), and Falcon  (Almazrouei et al., 2023), can also create higher extraction risks.
In this paper, we study a more realistic scenario and explore the actual risks posed by extraction attacks. More specifically, we ask:
@@
    Can adversaries exploit prompt sensitivity? We find that extraction attacks are sensitive to the prompt design, extracting over $20\%$ more data with even minor, unintuitive changes to the prompt (§5.1). Thus, an adversary, given the opportunity to prompt the model multiple times, can extract more data than previously observed.
@@
    Does access to multiple checkpoints increase extraction risk? An adversary with access to multiple model checkpoints over time or sizes gains broader access to the underlying dataset. We find such an adversary can increase the extraction rates up to 1.5×1.5$\times$1.5 ×, significantly heightening the risk of information leakage (§5.2).
@@
    Is data deduplication effective in reducing the extraction risks? We find that data deduplication does reduce the extraction risks, in line with the existing literature (Carlini et al., 2023). However, adversaries can still exploit the prompt structure and multiple checkpoints to extract more information (§6.3). Thus, our concerns about a powerful real-world adversary persist despite deduplication.
@@
    How are downstream applications affected by the presence of such an adversary? We performed three separate case studies and found that our more realistic adversary improves the $p$-value of dataset inference up to 2×2$\times$2 ×7.1), the extraction of copyright violations by up to $20\%$7.2), and the extraction rate of personally identifiable information (PIIs) by 1.5×1.5$\times$1.5 ×7.3).
@@ 2 Background and Related Work
In this section, we introduce relevant background on extraction attacks in LLMs, followed by an overview of related work on prompt sensitivity and training dynamics in LLMs. Finally, we describe the term churn as it applies in our context.
@@
Extraction Attacks in LLMs. Unintended memorization in LLMs can make it prone to information leakage (Tirumala et al., 2022; Carlini et al., 2019; Mattern et al., 2023; Carlini et al., 2022), particularly through extraction attacks (Birch et al., 2023; Carlini et al., 2021, 2023; Nasr et al., 2023). These attacks allow adversaries to extract training data from the model, raising concerns of leaking sensitive information (Birch et al., 2023). Extraction attacks have gained significant attention in recent years, studied under two primary frameworks: Discoverable Memorization (Carlini et al., 2023; Kassem et al., 2024; Liu et al., 2023b; Biderman et al., 2023a; Tirumala et al., 2022; Huang et al., 2022), where the adversary attempts to extract targeted information, and Extractable Memorization Nasr et al. (2023); Kandpal et al. (2022); Qi et al. (2024), where the adversary attempts to extract any information about the data.
@@
We add to the growing body of research on targeted extraction attacks by highlighting the lack of a realistic adversary in the literature. We show the existence of a stronger real-world adversary capable of combining information from various attacks, thereby defining a composite form of discoverable memorization (§3). Schwarzschild et al. (2024) also redefines discoverable memorization, using prompt optimization and adversarial compression ratio (ACR) to quantify memorization as information compression. In contrast, rather than relying on optimization, our focus is instead on exploiting the multi-faceted access to LLM training data.
@@
Prompt Sensitivity in LLMs. LLMs are shown to be sensitive to changes in their prompts, leading to fluctuations in their performance (Sclar et al., 2024; Liu et al., 2023a). This sensitivity persists across varying model sizes and through fine-tuning and other downstream modifications (Salinas and Morstatter, 2024; Zhu et al., 2023). The sensitivity of prompts can also be misused, and adversarial modifications to prompts can trigger the model to act in unintended ways Rossi et al. (2024); Liu et al. (2024); Hubinger et al. (2024); Liu et al. (2024). While several overarching trends studying the impact of prompt design on extraction attacks are present in the literature (Carlini et al., 2023; Kassem et al., 2024; Qi et al., 2024; Tirumala et al., 2022), these trends are often evaluated in isolation. Motivated by the composability of privacy leakage (McSherry, 2009), we argue that an adversary capable of repeated prompting can combine these trends. We show that such an adversary can extract more information about the training data than previously reported in the literature (§5.1).
@@
Training Dynamics of LLMs. Several recent works have studied the training dynamics of LLMs over time (Tirumala et al., 2022; Liu et al., 2021; Xia et al., 2023). In the context of memorization, recent work by Biderman et al. (2023a) explored the impact of model size and intermediate checkpoints on the dynamics of memorization, revealing a considerable variance in memorized data over time and size. The practice of releasing models in various sizes and regularly updating them over time can thus increase the attack surface for the underlying dataset. In our work, we study how adversaries can exploit access to multiple checkpoints of a model to extract more data (§5.2).
@@
Churn. The instability of model predictions under updates has gained significant attention in recent years, studied under the umbrella of churn (Milani Fard et al., 2016; Cotter et al., 2019; Bahri and Jiang, 2021; Anil et al., 2018; Jiang et al., 2021; Adam et al., 2023; Watson-Daniels et al., 2024). Churn quantifies the inconsistency in predictions between a system pre-update vs post-update, by measuring the fraction of examples whose predictions diverge (Milani Fard et al., 2016). The term churn is traditionally used in the literature to describe such regressive trends in model predictions, we extend its use by highlighting similar regressive trends and instability of extraction attacks under changing prompts and models. Thus, churn occurs when information is extractable with weaker setups like shorter prompts, smaller models, or earlier checkpoints, but not with the stronger setup.
@@ 3 Re-evaluating Adversarial Strengths
The adversary is central to our work. We begin by defining its capabilities, arguing that existing work has underestimated the strength of real-world adversaries. To ensure broad applicability, we assume gray-box access to the target model, i.e., the adversary can only access the generation output and probabilities from the model. Consequently, they cannot access model weights, gradients, or even control the generation hyperparameters, which reflects the typical level of accessibility for most commercial LLMs. Despite these constraints, we will demonstrate that an adversary in the current LLM ecosystem possesses far greater power than what has been recognized in existing literature.
@@ 3.1 Adversary Capabilities
Composability (or self-composability) of privacy leakage (McSherry, 2009) suggests that when an adversary gains access to multiple outputs from algorithms on the same underlying dataset—whether through multiple queries from the same algorithm or queries across multiple algorithms—the risk of information leakage grows. Consequently, an adversary with multiple points of access is significantly stronger than one with only a single point of access. In the current landscape of LLMs, such access is not only unsurprising but also easily obtainable (as illustrated in Figure 1). Specifically, we consider two forms of multi-faceted access:
@@
Exploiting Prompt Sensitivity LLMs are highly sensitive to their input, including its structure, content, and even the presence of noisy text within the prompt (Sclar et al., 2024; Liu et al., 2023a; Salinas and Morstatter, 2024; Zhu et al., 2023). While existing studies have focused on improving the prompts for stronger attacks, the nuance of prompt sensitivity in LLMs often defies intuitive expectations. For instance, while longer prompts are known to increase the success of extraction attacks Carlini et al. (2023), our work demonstrates that even shorter prompts can at times exploit vulnerabilities that longer prompts overlook (§5.1).
@@
Given the widespread use of LLMs through both chat interfaces and API calls, restricting model access is not realistic. While most commercial LLMs do have rate limits, they are quite high to be of practical concern. For example, even at the lowest tier subscription of \$$\$5$, ChatGPT has a 500 query per minute (qpm) rate limit for GPT4 and 3500 qpm for GPT3.5333qpm stats and subscription rate as of September 2024.. Thus, an adversary can prompt millions of generations in just one day, making it easier to exploit structural changes in prompts.
@@
Multiple Checkpoints. LLMs are typically deployed in various sizes to cater different needs for accuracy and efficiency among users. However, due to the stochastic nature of their training and the impact of scaling, different model sizes might memorize unique portions of the underlying dataset (Biderman et al., 2023a). Consequently, an adversary with access to multiple model sizes can effectively aggregate extracted information rather than limiting it to a single model (§5.2).
Similarly, deployed LLMs undergo regular updates driven by new data, better learning techniques, evolving security measures, and novel functionalities. The stochastic training process means that data resilient to attacks at a certain time step may become vulnerable in subsequent model updates, or vice-versa (Biderman et al., 2023a). Such fluctuations can enable adversaries to exploit multiple checkpoints over time, potentially extracting more information than from a static model (§5.2).
@@
More broadly, access to multiple models sharing common training data increases the attack surface, and in turn, creates stronger adversaries. This level of access is not unprecedented, and several companies in the current LLM ecosystem release multiple versions of their models and even update them periodically. For example, there are 8 different major versions of the ChatGPT models and more than 10 major versions of the Llama models currently available, while these models are also known to get regular minor updates accessible using update dates OpenAI (2024); Chen et al. (2023). Thus, access to multiple models trained on the same data, as has become commonplace, can significantly increase the risks of information leakage.
@@ 3.2 Combining Extraction Attacks
We argued for the heightened risk posed by multifaceted access to LLMs, either through repeated prompting or multiple model checkpoints. Before discussing our empirical study, we first quantify the risks associated with this stronger adversary. We argue that when an adversary gains such extensive access, any successful extraction of information—even if achieved once—renders that specific information vulnerable to the adversary.
Formally, adapting the definition of discoverable memorization from Nasr et al. (2023), we propose:
@@
For a set of $k$ models $\mathbb{G}=\{Gen_{i}|i\in[1\dots k]\}$, a set of $r$ prompt modifiers $\mathbb{F}=\{F_{j}|j=[1\dots r]\}$, and an example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ from the training set $\mathbb{X}$, we say ${\mathbf{x}}$ is composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}\;\textrm{s.t.}\;Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}$.
$CDM(\mathbb{G},\mathbb{F},$ ${\mathbf{p}}\parallel{\mathbf{x}})$
$=\max_{Gen_{i}\in\mathbb{G},F_{j}\in\mathbb{F}}\mathbbm{1}_{Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}}$
@@
Extraction attacks are often evaluated in the literature using a verbatim match  (Carlini et al., 2021, 2023; Nasr et al., 2023; Huang et al., 2022), i.e., the generated text must match the original text perfectly. However, this rigid metric does not take into account the noise in LLM generations, and several recent works have turned to approximate matching to quantify extraction risks for LLMs (Qi et al., 2024; Kassem et al., 2024; Liu et al., 2023b; Ippolito et al., 2022). Thus, we also extend our definition of composite extraction attacks to the approximate matching setup:
@@
For a set of $k$ models $\mathbb{G}=\{Gen_{i}|i\in[1\dots k]\}$, a set of $r$ prompt modifiers $\mathbb{F}=\{F_{j}|j=[1\dots r]\}$, a similarity metric $S$, a similarity threshold $\delta$, and an example $[{\mathbf{p}}\parallel{\mathbf{x}}]$ from the training set $\mathbb{X}$, we say ${\mathbf{x}}$ is approximate composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}\;\textrm{s.t.}\;S\left(Gen_{i}\left(F_{j}\left(\mathbf{p}\right)\right),\mathbf{x}\right)\geq\delta$.
@@
Here, $S$ is a similarity metric defined as a function $S:(\mathcal{W}^{*}\times\mathcal{W}^{*})\rightarrow[0,1]$ that takes as input two strings $a,b\in\mathcal{W}^{*}$, and returns a score between 00 and $1$ to represent the similarity between the two input strings, and $\delta$ is a threshold that controls the degree of approximate matching.

### v3

@@ Abstract
Language models are prone to memorizing their training data, making them vulnerable to extraction attacks. While existing research often examines isolated setups, such as a single model or a fixed prompt, real-world adversaries have a considerably larger attack surface due to access to models across various sizes and checkpoints, and repeated prompting. In this paper, we revisit extraction attacks from an adversarial perspective—with multi-faceted access to the underlying data. We find significant churn in extraction trends, i.e., even unintuitive changes to the prompt, or targeting smaller models and earlier checkpoints, can extract distinct information. By combining multiple attacks, our adversary doubles ($2\times$) the extraction risks, persisting even under mitigation strategies like data deduplication. We conclude with four case studies, including detecting pre-training data, copyright violations, extracting personally identifiable information, and attacking closed-source models, showing how our more realistic adversary can outperform existing adversaries in the literature.
@@ 1 Introduction
Large language models (LLMs) have grown considerably in size (Meta AI, 2024; Zhao et al., 2023), and have become integral to a wide range of tasks such as knowledge retrieval, question answering, code generation, and machine translation.
To complement this growing scale, LLMs are often trained on large amounts of data (Penedo et al., 2024; Soboleva et al., 2023; Gao et al., 2020; Raffel et al., 2020) that may include private information, especially if scraped from the web. As LLMs are prone to memorizing the data they have been trained on, they can be prompted to expose sensitive contexts—making it easier for an adversary to extract information. Naturally, a question arises: how big is the risk imposed due to memorization?
Extraction attacks offer an empirical framework to quantify the information leakage in the presence of an adversary. A commonly studied extraction attack is discoverable memorization (Carlini et al., 2023; Kassem et al., 2024), where the adversary extracts targeted information from the training data by prompting the model with a portion of a sentence from the training data to extract the rest. Discoverable memorization has been used in many adversarial settings, including membership inference (Maini et al., 2024), data contamination detection (Ravaut et al., 2024), and copyright violations (Karamolegkou et al., 2023), among others.
Current extraction attacks study memorization trends in LLMs across isolated settings like model sizes, generation hyperparameters, and learning dynamics (Carlini et al., 2023). While effective, they underestimate the risk posed due to a multi-faceted access to the underlying data in the current LLM ecosystem. For instance, we show that an adversary can exploit the sensitivity of LLMs to prompt structure, length, and content, to amplify the information gained. The current accessibility to frequently updated model sizes (Meta AI, 2024); checkpoints (Biderman et al., 2023b; Groeneveld et al., 2024) and a large array of model families such as Llama (Meta AI, 2024), Gemini (Team et al., 2023), and Falcon (Almazrouei et al., 2023), can also create higher extraction risks.
In this paper, we study a more realistic scenario and explore the actual risks posed by composite extraction attacks that can combine information from multiple attacks. More specifically, we ask:
@@
    Can adversaries exploit prompt sensitivity and access to multiple checkpoints? We find that extraction attacks are sensitive to the prompt design, extracting over $20\%$ more data with even minor, unintuitive changes (§5.1). Similarly, we find that an adversary with access to multiple model checkpoints can increase the extraction rates up to $1.5\times$5.2). Thus, an adversary with multi-faceted access can extract far more data than previously observed.
@@
    Should the adversary always attack the most vulnerable setup? Vulnerable setups are also more expensive to attack. Interestingly, we find that on a limited adversarial budget, using composite attacks on less vulnerable but cheaper setups can cause more information leakage (§6.2).
@@
    Is data deduplication effective against composite attacks? We find that data deduplication does reduce memorization, as expected (Carlini et al., 2023). However, adversaries can still exploit the prompt structure and multiple checkpoints to extract more information (§6.4), hence our concerns persist despite deduplication.
@@
    How are downstream applications affected by a stronger adversary? We performed four separate case studies and found that our more realistic adversary improves the $p$-value of dataset inference in open-source models by up to $2\times$7.1), the extraction of copyright violations by up to $20\%$7.2), the extraction rate of personally identifiable information (PIIs) by $1.5\times$7.3), and training data inference even in closed-source models by $16\%$7.4).
@@ 2 Background and Related Work
We first introduce relevant background on extraction attacks in LLMs, followed by an overview of related work on prompt sensitivity and training dynamics in LLMs. Finally, we describe the term churn as it applies in our context.
@@
Extraction Attacks in LLMs. Unintended memorization in LLMs can make it prone to information leakage (Tirumala et al., 2022; Carlini et al., 2019; Mattern et al., 2023; Carlini et al., 2022), particularly through extraction attacks (Birch et al., 2023; Carlini et al., 2021, 2023; Nasr et al., 2023). Extraction attacks have gained significant attention in recent years, studied under two primary frameworks: Discoverable Memorization (Carlini et al., 2023; Kassem et al., 2024; Liu et al., 2023b; Biderman et al., 2023a; Tirumala et al., 2022; Huang et al., 2022), where the adversary attempts to extract targeted information, and Extractable Memorization Nasr et al. (2023); Kandpal et al. (2022); Qi et al. (2024), where the adversary attempts to extract any information about the training data.
@@
We add to the growing body of research on targeted extraction attacks by highlighting the lack of a realistic adversary in the literature. We show the existence of a stronger adversary capable of combining information from various attacks, thus defining composite discoverable memorization3).
@@
Prompt Sensitivity in LLMs. LLMs are sensitive to changes in their prompts, leading to fluctuations in their performance (Sclar et al., 2024; Liu et al., 2023a). This sensitivity persists across varying model sizes and through fine-tuning and other downstream modifications (Salinas and Morstatter, 2024; Zhu et al., 2023). The sensitivity of prompts can also be misused, and adversarial modifications to prompts can trigger the model to act in unintended ways Rossi et al. (2024); Liu et al. (2024); Hubinger et al. (2024); Liu et al. (2024).
While several overarching trends studying the impact of prompt design on extraction attacks are present in the literature (Carlini et al., 2023; Kassem et al., 2024; Qi et al., 2024; Tirumala et al., 2022), these trends are often studied in isolation. Motivated by the composability of privacy leakage (McSherry, 2009), we argue that an adversary capable of repeated prompting can combine these trends, and extract more information about the training data than previously reported (§5.1).
@@
Training Dynamics of LLMs. Several recent works have studied the training dynamics of LLMs over time (Tirumala et al., 2022; Liu et al., 2021; Xia et al., 2023). In the context of memorization, Biderman et al. (2023a) explored the impact of model size and intermediate checkpoints on the dynamics of memorization, revealing a considerable variance in memorized data over time and size. The practice of releasing models in various sizes and regularly updating them can thus increase the attack surface. Similar to prompt sensitivity, we study how adversaries can also exploit access to multiple checkpoints to extract more data (§5.2).
@@
Churn. Churn quantifies the instability of model predictions under updates (Milani Fard et al., 2016; Bahri and Jiang, 2021; Jiang et al., 2021; Adam et al., 2023; Watson-Daniels et al., 2024), i.e., the inconsistency in predictions between a system pre-update vs post-update, by measuring the fraction of examples whose predictions diverge (Milani Fard et al., 2016). While the term churn is traditionally used to describe regressive trends in model predictions, we extend it by highlighting similar regressive trends of extraction attacks under changing prompts and models. Thus, churn occurs when information not extractable with a stronger setup is instead extractable with weaker setups like shorter prompts, smaller models, or earlier checkpoints.
@@ 3 Re-evaluating Adversarial Strengths
The adversary is central to our work. We start by defining its capabilities and argue that prior work underestimates real-world adversaries. To ensure broad applicability, we assume gray-box access: the adversary can observe model outputs and probabilities but cannot access weights, gradients, or even control generation hyperparameters, typical in commercial LLMs. Despite these constraints, we show that adversaries are far more powerful than previously recognized due to their multi-faceted access to the underlying data (see Figure 1).
We focus on discoverable memorization, i.e., we assume access to the ground-truth completion to test whether the extracted information is correct. Here, the adversary is primarily interested in auditing the memorization behaviour of the model. This is central to many applications, including membership inference, dataset inference, copyright violations, privacy auditing, among others, revisited in §7. That said, as we argue in §8, the larger attack surface and implications of a stronger adversary remain even beyond discoverable memorization.
@@ 3.1 Adversary Capabilities
Composability (or self-composability) in privacy (McSherry, 2009) implies that access to multiple outputs from the same data increases the risk of information leakage. Thus, an adversary with multiple access points is much more powerful. In the current landscape of LLMs, such access is not only unsurprising but easily obtainable. Specifically, we focus on two forms of multi-faceted access:
@@
Exploiting Prompt Sensitivity. LLMs are highly sensitive to their input, including its structure and content (Sclar et al., 2024; Liu et al., 2023a; Salinas and Morstatter, 2024; Zhu et al., 2023). While existing studies have focused on improving the prompts for stronger attacks, the nuance of prompt sensitivity in LLMs often defies intuitive expectations. For instance, while longer prompts are known to increase the success of extraction attacks Carlini et al. (2023), our work demonstrates that even shorter prompts can at times exploit vulnerabilities that longer prompts overlook (§5.1).
@@
Given the widespread use of LLMs through both chat interfaces and API calls, restricting model access is not realistic. While most commercial LLMs do have rate limits, they are quite high to be of practical concern. For example, even at the lowest tier subscription of \$5\\$5, ChatGPT has a 500 query per minute (qpm) rate limit for GPT4 and 3500 qpm for GPT3.5111qpm stats and subscription rate as of September 2024.. Thus, an adversary can prompt millions of generations in just one day, making it easier to exploit prompt sensitivity.
Figure 2: A toy illustration of how churn may emerge across completions from different model sizes. Adversaries can utilize this churn to increase the number of valid extractions.
@@
Multiple Checkpoints. LLMs are typically deployed in various sizes to cater to different needs for accuracy and efficiency among users. They also undergo regular updates driven by new data, better learning techniques, evolving security measures, and novel functionalities. Due to the stochastic nature of their training and the impact of scaling, different model sizes or checkpoints might memorize unique portions of the underlying data (Biderman et al., 2023a). Consequently, an adversary with access to various checkpoints across sizes or training can aggregate extracted information (§5.2).
@@
Thus, access to multiple models trained on overlapping datasets substantially increases the attack surface, thereby amplifying the capabilities of adversaries. This level of access has become increasingly common in the current LLM ecosystem. For example, there exist over eight major versions of ChatGPT and ten major versions of Llama, alongside regular minor updates OpenAI (2024); Chen et al. (2023). As such, the availability of multiple models with shared training data can significantly increase the risks of information leakage.
@@ 3.2 Combining Extraction Attacks
We discussed the elevated risks of multi-faceted access to the training data. Before presenting our empirical results, we first quantify the risks associated with this stronger real-world adversary. We argue that once the adversary gains such broad access, any successful extraction—even if achieved once—renders that specific information vulnerable to the adversary (see Figure 2).
Formally, for a sentence ${\mathbf{p}}\parallel{\mathbf{x}}]$ in the training data, where $\parallel$ represents the concatenation of two strings, the adversary has access to the prompt $p$ and aims to extract information $x$. Adapting the definition of discoverable memorization from Nasr et al. (2023), we propose:
@@
For a set of models $\mathbb{G}=\{{Gen_{i}|}_{i=1}^{k}\}$, prompt modifiers $\mathbb{F}=\{{F_{j}|}_{j=1}^{r}\}$, and an example ${\mathbf{p}}\parallel{\mathbf{x}}]$ from the training dataset $\mathbb{X}$, we say $\mathbf{x}$} is composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}$, such that, $Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}$.
$CDM(\mathbb{G},\mathbb{F},$ ${\mathbf{p}}\parallel{\mathbf{x}})=\max_{Gen_{i}\in\mathbb{G},F_{j}\in\mathbb{F}}\mathbbm{1}_{Gen_{i}(F_{j}(\mathbf{p}))=\mathbf{x}}$
@@
Extraction attacks are often evaluated in the literature using a verbatim match  (Carlini et al., 2021, 2023; Nasr et al., 2023; Huang et al., 2022), i.e., the generated text must match the original text perfectly. However, this rigid metric does not take into account the noise in LLM generations, and several works have turned to approximate matching (Qi et al., 2024; Kassem et al., 2024; Liu et al., 2023b; Ippolito et al., 2022). Thus, we also extend our definition of composite extraction attacks to the approximate matching setup as:
@@
For a set of models $\mathbb{G}=\{{Gen_{i}|}_{i=1}^{k}\}$, prompt modifiers $\mathbb{F}=\{{F_{j}|}_{j=1}^{r}\}$, a similarity metric and threshold $S,\delta$, and an example ${\mathbf{p}}\parallel{\mathbf{x}}]$ from the training dataset $\mathbb{X}$, we say $\mathbf{x}$} is approximate composite discoverably memorized if $\exists\;Gen_{i}\in\mathbb{G}\;\textrm{and}\;F_{j}\in\mathbb{F}$, such that, $S\left(Gen_{i}\left(F_{j}\left(\mathbf{p}\right)\right),\mathbf{x}\right)\geq\delta$.
@@
Here, $S$ is a similarity metric defined as a function $S:(\mathcal{W}^{*}\times\mathcal{W}^{*})\rightarrow[0,1]$ that takes as input two strings $a,b\in\mathcal{W}^{*}$, and returns a score between 0 and $1$ to represent the similarity between the two input strings, and $\delta$ is a threshold that controls the degree of approximate matching. We experiment with various similarity metrics $S$ in §6.3.